OpenStudy (anonymous):

Simplifying Rational Expressions, I don't understand how they got this answer?

OpenStudy (anonymous):

\[2p-24\div 4p-48\] This is the original problem.

OpenStudy (anonymous):

First I canceled out the 24 and divided 48 by 24 to get 2. then i cancelled out 2p and divided 4p and got 2.

OpenStudy (anonymous):

\[1-1/2-2\] This is the answer I got.

OpenStudy (anonymous):

The book has 1/2 written as the answer.I don't understand that.

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\frac{2p-24}{4p-48}\]

OpenStudy (mertsj):

First you must know that only factors may be cancelled. That's why you are having trouble. You cancelled addends.

OpenStudy (mertsj):

So before you can cancel, you have to factor the numerator and denominator.

OpenStudy (mertsj):

\[\frac{2x-24}{4x-48}=\frac{2(x-12)}{4(x-12)}\]

OpenStudy (mertsj):

Now...those are factors and can be cancelled. Can you finish it?

OpenStudy (anonymous):

Ok, I get that :) Thank you for showing me :)

OpenStudy (mertsj):

yw